# Topological flow data analysis for transient flow patterns: a graph-based approach

### v1

@@ Abstract.
We introduce a time-series analysis method for transient two-dimensional flow patterns based on Topological Flow Data Analysis (TFDA), a new approach to topological data analysis. TFDA identifies local topological flow structures from an instantaneous streamline pattern and describes their global connections as a unique planar tree and its string representation. With TFDA, the evolution of two-dimensional flow patterns is reduced to a discrete dynamical system represented as a transition graph between topologically equivalent streamline patterns. We apply this method to study the lid-driven cavity flow at Reynolds numbers ranging from $Re=14000$ to $Re=16000$, a benchmark problem in fluid dynamics data analysis. Our approach reveals the transition from periodic to chaotic flow at a critical Reynolds number when the reduced dynamical system is modelled as a Markov process on the transition graph. Additionally, we perform an observational causal inference to analyse changes in local flow patterns at the cavity corners and discuss differences with a standard interventional sensitivity analysis. This work demonstrates the potential of TFDA-based time-series analysis for uncovering complex dynamical behaviours in fluid flow data.
@@ 1. Introduction
Investigations of fluid dynamics are not complete when we finish collecting data on the flow field, such as the velocity, pressure, vorticity and so forth. To describe and understand the rich behaviours of a flow in a given condition, we need to find a structure in the data. Some structures are easily caught by our eyes, like a vortex sheet or tube, which can be highlighted by certain visualization methods in laboratory experiments or numerical simulations. A hairpin vortex in the boundary layer turbulence is a primary example of those structures that lead to the current understanding of the turbulence, see e.g., Davidson (2015). However, finding a relevant structure of a given flow is hard in general. Moreover finding a reduced-order model of the flow with respect to the elements of the structure is even harder. For the reduced-order modelling, the Galerkin projection with handful terms of the known governing equation has been a systematic tool. The data-driven approaches without needing to know the governing equation include so-called proper orthogonal decomposition (POD) and dynamic mode decomposition (DMD), as explained in Brunton and Kutz (2022) for instance.
@@
Topological Data Analysis (TDA) is a relatively new method describing the topological features of data. See an introductory book by Edelsbrunner and Harer (2010) and an advanced book by Dey and Wang (2022). While TDA applications in fluid dynamics have gained attention in recent years, concrete examples of its implementation remain limited. In the present paper, we construct a reduced model of the temporal evolution of streamline patterns using a different TDA approach, called Topological Flow Data Analysis (TFDA), that focuses on the topology of instantaneous streamline patterns of two-dimensional flows. Through this model, we quantitatively track the temporal changes in the topological features of the flow field and analyse the causal relationships between local flow structures by treating pattern transitions as a discrete dynamical system on the reduced model. We thus discuss both the effectiveness of this method and its potential for future development.
@@
The mathematical background behind the TFDA is explained as follows. Let $\psi({\bm{x}})$ be a streamfunction of $C^{r}$-class ($r\geq 1$) on a two-dimensional domain $\mathcal{D}\subset\mathbb{R}^{2}$. Since the vector field is defined by ${\bm{u}}=-\nabla^{\perp}\psi$ with $\nabla^{\perp}=(\partial_{y},-\partial_{x})$, it is regarded as a Hamiltonian vector field with $\psi$ being its Hamiltonian function. Let $O({\bm{x}}_{0})$ denote a particle orbit advected by the vector field ${\bm{u}}({\bm{x}})$ starting from ${\bm{x}}_{0}\in\mathcal{D}$. That is to say, $O({\bm{x}}_{0})=\{{\bm{\phi}}(s;{\bm{x}}_{0})\,|\,s\in\mathbb{R}\}$, where ${\bm{\phi}}(s;{\bm{x}}_{0})$ is the unique solution to the initial value problem $\dot{\bm{\phi}}(s)={\bm{u}}({\bm{\phi}}(s))$ and ${\bm{\phi}}(0)={\bm{x}}_{0}$. When the vector field is stationary, the particle orbit is contained in a level curve of the streamfunction $\psi$ owing to $\nabla\psi\cdot{\bm{u}}=0$. A mathematical theory for classifying topological features of the set of all particle orbits, $\mathcal{O}=\{O({\bm{x}}_{0})\,|\,{\bm{x}}_{0}\in\mathcal{D}\}$, have been developed by Yokoyama and Sakajo (2013); Sakajo and Yokoyama (2018). The topological classification is realised under the restrictions of Poincaré–Bendixson theorem, which gives a classification of fixed points, periodic orbits, and limit orbits as $s\rightarrow\pm\infty$, and Poincaré–Hopf theorem, which defines the relationship between the index of the vector field and the Euler characteristic of the flow domain. Based on the theory, it is shown in Uda et al. (2019) that the topological structure of particle orbits in $\mathcal{O}$ is converted uniquely into a discrete graph, named a partially Cyclically Ordered rooted Tree (the abbreviation COT is used in this paper), and its string expression, called COT representation. Through this transformation, the topological patterns of flow streamlines are represented as graphs or strings, which we call Topological Flow Data Analysis (TFDA). It can also describe the evolution of complex flow patterns as a discrete dynamical system of trees, which gives rise to a reduced-order model of complex flow evolution in terms of topology, as we will demonstrate in this paper. Note that TFDA was originally developed for Hamiltonian vector fields as above, but it has been extended to two-dimensional compressible vector fields recently (Sakajo and Yokoyama, 2022).
@@
TFDA has been applied to a variety of fluid mechanical problems. In Sakajo et al. (2014), TFDA was used to analyse boundary-layer separations around a flat plate with the chord-length based Reynolds number $1370$ and the angle of attack $15^{\circ}$ in a numerical simulation of a two-dimensional flow. It was shown that the dynamical process was described with brief COT representations. In Uda et al. (2021), TFDA was employed to detect an atmospheric blocking phenomenon, a long-lived high-pressure region blocking the westerly jet flow in the surrounding region, with a year-long observational data in the synoptic scale (over $1000$ km). It was confirmed that the method of TFDA outperformed other existing methods for detecting the blocking and that morphological types of the blocking were identified with the COT representations in accordance with the meteorologists’ empirical classifications. Sakajo et al. (2022) considered Kuroshio Large Meander in oceanography, which is an abnormal current that gives rise to the entrapment of a cool cyclonic eddy that has substantial impacts on fishery and marine transport. TFDA is applied to sea surface height data to detect the occurrence of the abnormal state. Furthermore, in Sakajo and Itatani (2023), TFDA has been utilised to classify flow patterns in the left ventricle of the human heart. Actual measurement data with an echocardiography (two spacial dimensions and time) and magnetic resonance imaging (a slice of three spacial dimensions and time) were analysed. Specifically, one can define a swirling flow domain as a topological vortex structure using the COT representations, and it was demonstrated that differences in the flow patterns between normal and diseased hearts were resolved objectively with topological vortex structures in agreement with diagnoses of experienced cardiovascular surgeons.
@@
In these applications mentioned above, TFDA was mainly used to state identifications of flow patterns with vortex-like flow structures, but methods for characterizing the dynamics of transient flow patterns using TFDA have not yet been fully developed. Therefore, in this paper, we propose a new time-series analysis based on TFDA to describe the complex dynamics exhibited by transient flow patterns using a graph-theoretic approach. Our target flow is simple and prototypical: the two-dimensional lid-driven cavity flow (Shen, 1991). It is perhaps the most famous benchmark problem in computational fluid mechanics when the flow is in a nonlinear steady-state consisting of several corner eddies and one big central eddy. The flow is also regarded as a representative example of the complexity of recirculating eddy and flow separation in a textbook (Ruban and Gajjar, 2013). Those eddies are explicitly expressed in the COT representation, as we will see later. This means that the tree representation obtained via TFDA is always interpretable in contrast to, for example, the singular vectors of POD and DMD, which are not always easy to interpret. In particular, we demonstrate the capability of TFDA by applying it to the numerical lid-driven cavity flow with the transitional range of Reynolds numbers from time-periodic states to chaotic states.
The paper is constructed as follows. In Section 2, after reviewing the lid-driven cavity flow and the numerical methods to compute it, we provide how we apply TFDA to the lid-driven cavity flows. Section 3 provides numerical results of the time-series analysis proposed in this paper. In Section 3.1, we first describe how snapshots of streamline patterns in the lid-driven cavity flow are converted into COT and COT representations at a moderate Reynolds number. We then show that the evolution of the cavity flow is expressed as a hopping transition among a couple of the COT representations, which is visualised as a network among the COT representations as nodes. This leads to a discrete reduced-order dynamical model of the lid-driven cavity flow. In Section 3.2, we demonstrate that the transition from the periodic state to the chaotic state over Reynolds numbers corresponds to a complication of the network. This hopping picture yields a transition matrix among the tree representations by assuming that the hopping is Markovian. We then utilise a discrete graph analysis to transition matrices to characterize the evolution of flow patterns. Furthermore, to take advantage of the high interpretability of TFDA, we perform an observational causal analysis based on the time series of the tree representation, which is known as the convergent cross mapping (CCM) developed by Sugihara et al. (2012). Namely, we address which corner eddy is more causal than the other one. We show that the observational CCM approach based on TFDA detects a geometric causality between topological changes in the chaotic flow patterns that the standard sensitivity analysis cannot capture as shown in Appendix A. Discussion and concluding remarks are in Section 4.
@@ 2. Our target and method
2.1. Physics of the lid-driven cavity flow
The lid-driven cavity flow is a viscous flow in a hollow space enclosed by no-slip walls. The shape of the cavity is usually set to a rectangle in two spatial dimensions and to a rectangular cuboid in three dimensions. One of the walls, the lid, moves with a constant velocity and other walls stay still (zero velocity) as shown in Figure 1(a). As the energy constantly injected by the lid is dissipated by the viscosity, the flow reaches a variety of states depending on the Reynolds number as reviewed in Shankar and Deshpande (2000). For low Reynolds numbers, the flows are stationary. In other words, they are non-linearly stationary solutions to the Navier-Stokes equations with the given boundary condition. Those stationary solutions, in particular two-dimensional ones, serve as a standard benchmark for numerical methods, see e.g., Ghia et al. (1982) and Botella and Peyret (1998). As we increase the Reynolds number, the flow becomes time-periodic and then turbulent (Shankar and Deshpande, 2000).
A fundamental fluid-dynamical characteristic of the lid-driven cavity flow, whether it is stationary, periodic or turbulent, is the formation of recirculating eddies at the corners, as emphasized by Ruban and Gajjar (2013). This phenomenon is illustrated in Figure 1(b), which shows the streamfunction of a stationary solution. In non-stationary flows, the separation and merger of these eddies and boundary layers become important. As Shankar and Deshpande (2000) notes, the lid-driven cavity flow exhibits the full spectrum of fluid mechanical phenomena within the simplest geometrical configuration. This naturally raises the question: how can we systematically understand the flow variations and transitions through the dynamics of these prominent eddies? As we will show in this paper, TFDA provides a systematic and mathematically rigorous framework to achieve this goal for dynamic flow evolutions.
@@ 2.1. Physics of the lid-driven cavity flow
We consider a two-dimensional lid-driven cavity flow in a square cavity. The movement of the lid is regularised around the corners. We adopt the regularisation proposed in Shen (1991). Such a regularisation may not be necessary if the corner singularities are properly treated (Batchelor, 1967; Botella and Peyret, 1998; Shankar and Deshpande, 2000). Our adoption of regularisation is for the sake of numerical simplicity. We specifically consider the incompressible Navier–Stokes equations,
@@
in the square domain $\Omega=\{(x,y)\in\mathbb{R}^{2}|-1\leq x\leq 1,-1\leq y\leq 1\}$. Hence the side length of the square is $L=2$. Here ${\bm{u}}(x,y,t)$ is the velocity field, $p$ is the pressure and $\nu$ is the kinematic viscosity. The fluid density is normalised to one. The boundary condition is ${\bm{u}}={\bm{0}}$ on the boundary walls, except for the lid at the top. The velocity of the lid is given by
@@
where $u(x,y,t)$ and $v(x,y,t)$ are the $x$ and $y$-components of the velocity (Shen, 1991). Note that the velocity of the lid becomes zero smoothly at the corners ($x=\pm 1$). The Reynolds number here is defined with the maximum velocity of the lid, $U=\max_{-1\leq x\leq 1}u(x,y=1,t)=1$, and the side length of the square, $L=2$, as
@@
The numerical scheme we use is the Chebyshev-tau method in the streamfunction and vorticity formulation of the Navier-Stokes equations. We use the $3/2$ rule to remove the dealiasing error. The number of Gauss–Lobatto collocation points is $65^{2}$. The boundary condition is numerically satisfied by using the influence matrix method (Peyret, 2002). Because of the well-known loss of the rank of the influence matrix, we need to set the vorticity to zero at four collocation points on the boundary (in addition to the corner points). We follow the choice of the four points proposed in Ehrenstein and Peyret (1989), namely the nearest collocation points to the corner points on the lid and bottom wall. The time marching scheme is semi-implicit. Specifically, we use the second-order Adams-Bashforth method for the nonlinear term and the second-order backward differentiation for the viscous term, which is known as AB/BDI2 scheme (Peyret, 2002). The time-step value is $\Delta t=1.0\times 10^{-3}$.
We numerically simulate the flow for Reynolds numbers from $Re=14000$ to $Re=16000$ in steps of $50$. For each $Re$, we start the simulation with the zero velocity field and continue it up to time $T_{s}\gg 1$ when the flow reaches a statistically stationary state to avoid initial transients. Then we start to analyse the flow for $T_{s}\leq t\leq T_{s}+\tilde{t}$. Their values are $T_{s}=5000=2500(L/U)$ and $\tilde{t}$ varying from $200=100(L/U)$ to $10000=5000(L/U)$. In Table 1, we list numerical parameters for three representative Reynolds numbers. The convergence study is done for $Re=15000$ and $Re=16000$ by increasing the collocation points to $129^{2}$ or $257^{2}$ and halving the time step $\Delta t$ to check that the results do not change.
As a global fluid mechanical quantity, we consider the kinetic energy
@@
and use it as an indicator of the state of the flow as in Shen (1991). We show the variation of the kinetic energy in Figure 2 and its power spectral density (PSD) in Figure 3. The PSD for $Re=14000$ indicates that the flow is periodic. For $Re=15500$, Shen (1991) concluded that the flow is quasi-periodic. The PSD shown in Figure 3(b) is consistent with the conclusion. For $Re=16000$, the PSD becomes broad and the fine structures seen at $Re=15500$ are swallowed by the broad parts. This indicates that the flow is chaotic at $Re=16000$.
$Re$ collocation points $\Delta t$ State (Shen, 1991) State (present)

### v2

@@ Abstract.
We introduce a method of time series analysis for two-dimensional transient flow patterns based on Topological Flow Data Analysis (TFDA), a new approach to topological data analysis. TFDA identifies local topological flow structures from an instantaneous streamline pattern and describes their global connections as a unique planar tree and its string representation. With TFDA, the evolution of two-dimensional flow patterns is reduced to a discrete dynamical system represented as a transition graph between topologically equivalent streamline patterns. We apply this method to study the lid-driven cavity flow for Reynolds numbers from $Re=14000$ to $16000$, a benchmark problem in the analysis of fluid dynamics. Our approach can extract some physical information from the lid-driven cavity flow: transition of the flow from periodic to quasi-periodic and chaotic; estimation of the period of periodic dynamics; relation between variations in energy and enstrophy and topological changes in flow patterns; statistical properties of intricate flow evolution at higher Reynolds number. In addition, we perform an observational causal inference to analyse changes in local flow patterns in the cavity corner. This work demonstrates the potential of TFDA-based time series analysis to uncover complex dynamical behaviours in fluid flow data from a topological perspective.
@@ 1. Introduction
Investigations of fluid dynamics are not complete when we finish collecting data on the flow field, such as velocity, pressure, vorticity, and so forth. To describe and understand the rich behaviours of a flow in a given condition, we need to find a structure in the data. Some structures are easily caught by our eyes, such as a vortex sheet or tube, which can be highlighted by certain visualization methods in laboratory experiments or numerical simulations. A hairpin vortex in the boundary layer turbulence is a primary example of those structures that lead to the current understanding of turbulence, see, e.g., Davidson (2015). However, in general, finding a relevant structure for a given flow is difficult. Moreover, finding a reduced-order model of the flow with respect to the elements of the structure is even harder. For reduced-order modelling, the Galerkin projection with a handful of terms of the known governing equation has been a systematic tool. Data-driven approaches without needing to know the governing equation include the so-called proper orthogonal decomposition (POD) and dynamic mode decomposition (DMD), as explained in Brunton and Kutz (2022) for instance.
@@
Topological Data Analysis (TDA) is a relatively new method that describes the topological features of data. See an introductory book by Edelsbrunner and Harer (2010) and an advanced book by Dey and Wang (2022). While TDA applications in fluid dynamics have gained attention in recent years, concrete examples of their implementation remain limited. In the present paper, we construct a reduced model of the temporal evolution of streamline patterns using a different TDA approach, called Topological Flow Data Analysis (TFDA), which focuses on the topology of instantaneous streamline patterns of two-dimensional flows. Through this model, we quantitatively track the temporal changes in the topological features of the flow field and analyse the causal relationships between local flow structures by treating pattern transitions as a discrete dynamical system on the reduced model. We discuss both the effectiveness of this method and its potential for future development.
@@
The mathematical background behind the TFDA is explained as follows. Let $\psi({\bm{x}})$ be a streamfunction of $C^{r}$-class ($r\geq 1$) on a two-dimensional domain $\mathcal{D}\subset\mathbb{R}^{2}$. Since the vector field is defined by ${\bm{u}}=-\nabla^{\perp}\psi$ with $\nabla^{\perp}=(\partial_{y},-\partial_{x})$, it is considered a Hamiltonian vector field with $\psi$ being its Hamiltonian. Let $O({\bm{x}}_{0})$ denote a particle orbit advected by the vector field ${\bm{u}}({\bm{x}})$ starting from ${\bm{x}}_{0}\in\mathcal{D}$. That is, $O({\bm{x}}_{0})=\{{\bm{\phi}}(s;{\bm{x}}_{0})\,|\,s\in\mathbb{R}\}$, where ${\bm{\phi}}(s;{\bm{x}}_{0})$ is the unique solution to the initial value problem $\dot{\bm{\phi}}(s)={\bm{u}}({\bm{\phi}}(s))$ and ${\bm{\phi}}(0)={\bm{x}}_{0}$. When the vector field is stationary, the particle orbit is contained in a level curve of the streamfunction $\psi$, i.e., streamlines, owing to $\nabla\psi\cdot{\bm{u}}=0$. When the velocity field depends on time, particle orbits do not correspond to streamlines in general. However, we consider the topological structure of instantaneous streamlines every fixed time. A mathematical theory for classifying topological features of the set of all instantaneous streamlines, $\mathcal{O}=\{O({\bm{x}}_{0})\,|\,{\bm{x}}_{0}\in\mathcal{D}\}$, has been developed by Yokoyama and Sakajo (2013); Sakajo and Yokoyama (2018). The topological classification is realised under the restrictions of Poincaré–Bendixson theorem, which gives a classification of fixed points, periodic orbits, and limit orbits as $s\rightarrow\pm\infty$, and Poincaré–Hopf theorem, which defines the relationship between the index of the vector field and the Euler characteristic of the flow domain. Based on the theory, it is shown in Uda et al. (2019) that the topological structure of instantaneous streamlines in $\mathcal{O}$ is uniquely converted into a discrete graph, named a partially Cyclically Ordered rooted Tree (the abbreviation COT is used in this paper), and its string expression, called COT representation. Through this transformation, the topological patterns of flow streamlines are represented as graphs or strings, which we call Topological Flow Data Analysis (TFDA). It can also describe the evolution of complex flow patterns as a discrete dynamical system of trees, which gives rise to a reduced-order model of complex flow evolution in terms of topology, as we will demonstrate in this paper. Note that TFDA was originally developed for Hamiltonian vector fields as above, but it has recently been extended to two-dimensional compressible vector fields Sakajo and Yokoyama (2023).
@@
TFDA has been applied to a variety of fluid-mechanical problems. In Sakajo et al. (2014), TFDA was used to analyse boundary layer separations around a flat plate with the Reynolds number $1370$ based on chord length and angle of attack $15^{\circ}$ in a numerical simulation of a two-dimensional flow. The dynamical process was described with brief COT representations. In Uda et al. (2021), TFDA was used to detect an atmospheric blocking phenomenon, a long-lived high-pressure region that blocks the westerly jet flow in the surrounding region, with observational data of a year on the synoptic scale of more than $1000$ km. It was confirmed that the TFDA method outperformed other existing methods for detecting blocking and that the morphological types of blocking were identified with the COT representations in accordance with the meteorologists’ empirical classifications. Sakajo et al. (2022) considered Kuroshio Large Meander in oceanography, which is an abnormal current that causes the capture of a cool cyclonic eddy with substantial impacts on fisheries and marine transport. TFDA is applied to sea surface height data to detect the occurrence of an abnormal state. Recently, Kimura et al. (2026) has applied a TFDA to two dimensional turbulent flows in a doubly periodic box, allowing us to track coherent vortex structures statistically. Furthermore, TFDA has been utilised more complicated flow patterns in the left ventricle of the human heart Sakajo and Itatani (2023); Dyer et al. (2026). Actual measurement data with echocardiography (two spatial dimensions and time) and magnetic resonance imaging (a slice of three spatial dimensions and time) were analysed. Specifically, a swirling flow domain is defined as a topological vortex structure using COT representations, and it was shown that differences in flow patterns between normal and diseased hearts were resolved objectively with topological vortex structures in agreement with the diagnoses of experienced cardiovascular surgeons.
@@
In these applications mentioned above, TFDA was mainly used to state identifications of flow patterns with vortex-like flow structures, but methods for characterizing the dynamics of transient flow patterns using TFDA have not yet been fully developed. Therefore, in this paper, we propose a new time series analysis based on TFDA to describe the complex dynamics exhibited by transient flow patterns. Our target flow is simple and prototypical: the two-dimensional lid-driven cavity flow Shen (1991). It is perhaps the most famous benchmark problem in computational fluid mechanics when the flow is in a nonlinear steady state consisting of several corner eddies and one big central eddy. The flow is also considered a representative example of the complexity of the recirculating eddy and flow separation in a textbook Ruban and Gajjar (2013). Those eddy structures are explicitly expressed in the COT representation, as we see later. This means that the graph representation obtained via TFDA is always interpretable in contrast to, for example, the singular vectors of POD and DMD, which are not always easy to interpret. In particular, we demonstrate the capability of TFDA by applying it to numerical lid-driven cavity flow with the transitional range of Reynolds numbers from time-periodic states to chaotic states.
The paper is constructed as follows. In Section 2, after reviewing the lid-driven cavity flow and the numerical methods used here, we provide how we apply TFDA to the numerical data. Section 3 provides a detailed time series analysis using TFDA when the Reynolds number is $Re=14000$. We identify the change in topological flow patterns from the COT representations and show that flow dynamics is reduced to a transition graph between topologically equivalent flow patterns, called topological states. In Section 4, we investigate the physics of the lid-driven cavity flow across Reynolds numbers from $Re=14000$ to $16000$, where the flow dynamics changes from periodic to quasi-periodic and chaotic. We first show how the flow dynamics becomes complicated in terms of the number of topological states as the Reynolds number increases. Second, we estimate the period of periodic flow dynamics from the time series of COT representations. Third, we relate the variations of physical quantities such as energy and enstrophy to the transitions of topological states. Fourth, we assign serial numbers (IDs) to topological states in the order of their appearance across Reynolds numbers, and examine whether these states persist in the transition to chaotic flow through their occurrence rates. Finally, to take advantage of the high interpretability of TFDA, we perform an observational causal analysis based on the time series of COT representations.We show that the observational approach based on TFDA detects a geometric causality between local topological patterns in chaotic flow that the standard sensitivity analysis cannot capture, as compared in Appendix A. The discussion and concluding remarks are in Section 5.
@@ 2. Our target and method
2.1. The lid-driven cavity flow
The lid-driven cavity flow is a viscous flow in a hollow space enclosed by no-slip walls. The shape of the cavity is usually set to a rectangle in two spatial dimensions and to a rectangular cuboid in three dimensions. One of the walls, the lid, moves with a constant velocity and other walls remain still (zero velocity), as shown in Figure 1(a). As the energy injected constantly by the lid is dissipated by the viscosity, the flow reaches a variety of states depending on the Reynolds number as reviewed in Shankar and Deshpande (2000). For low Reynolds numbers, the flows are stationary. In other words, they are nonlinearly stationary solutions to the Navier-Stokes equations with the given boundary condition. Those stationary solutions, in particular two-dimensional ones, serve as a standard benchmark for numerical methods; see, e.g., Ghia et al. (1982) and Botella and Peyret (1998). As we increase the Reynolds number, the flow becomes time-periodic and then turbulent Shankar and Deshpande (2000).
A fundamental fluid-dynamical characteristic of the lid-driven cavity flow, whether stationary, periodic, or turbulent, is the formation of recirculating eddies at the corners, as emphasized by Ruban and Gajjar (2013). This phenomenon is illustrated in Figure 1(b), which shows the streamfunction of a stationary solution. In non-stationary flows, the separation and merger of these eddies and boundary layers become important. As Shankar and Deshpande (2000) notes, the lid-driven cavity flow exhibits the full spectrum of fluid mechanical phenomena within the simplest geometrical configuration. This naturally raises the question: How can we systematically understand the flow variations and transitions through the dynamics of these prominent eddies? As we show in this paper, TFDA provides a systematic and mathematically rigorous framework to achieve this goal for dynamic flow evolutions.
@@ 2.1. The lid-driven cavity flow
We consider a two-dimensional lid-driven cavity flow in a square cavity. The movement of the lid is regularised around the corners. We adopt the regularisation proposed in Shen (1991). Such a regularisation may not be necessary if the corner singularities are properly treated Batchelor (1967); Botella and Peyret (1998); Shankar and Deshpande (2000). Our adoption of regularisation is for the sake of numerical simplicity. We specifically consider the incompressible Navier–Stokes equations,
@@
in the square domain $\Omega=\{(x,y)\in\mathbb{R}^{2}|-1\leq x\leq 1,-1\leq y\leq 1\}$. Hence, the side length of the square is $L=2$. Here ${\bm{u}}(x,y,t)$ is the velocity field, $p$ is the pressure and $\nu$ is the kinematic viscosity. The fluid density is normalised to one. The boundary condition is ${\bm{u}}={\bm{0}}$ on the boundary walls, except for the lid at the top. The velocity of the lid is given by
@@
where $u(x,y,t)$ and $v(x,y,t)$ are the $x$ and $y$ components of the velocity Shen (1991). Note that the velocity of the lid smoothly decreases to zero at the corners ($x=\pm 1$). The Reynolds number here is defined with the maximum velocity of the lid, $U=\max_{-1\leq x\leq 1}u(x,y=1,t)=1$, and the side length of the square, $L=2$, as
@@
The numerical scheme we use is the Chebyshev-tau method in the streamfunction and vorticity formulation of the Navier-Stokes equations. We use the $3/2$ rule to remove the dealiasing error. The number of Gauss–Lobatto collocation points is $65^{2}$. The boundary condition is numerically satisfied by using the influence matrix method Peyret (2002). Because of the well-known loss of the rank of the influence matrix, we need to set the vorticity to zero at four collocation points on the boundary (in addition to the corner points). We follow the choice of the four points proposed in Ehrenstein and Peyret (1989), namely the nearest collocation points to the corner points on the lid and bottom wall. The time marching scheme is semi-implicit. Specifically, we use the second-order Adams-Bashforth method for the nonlinear term and the second-order backward differentiation for the viscous term, which is known as AB/BDI2 scheme Peyret (2002). The time-step value is $\Delta t=1.0\times 10^{-3}$.
We numerically simulate the flow for the Reynolds numbers from $Re=14000$ to $Re=16000$ in steps of $50$. For each $Re$, we start the simulation with the zero velocity field and continue until time $T_{s}\gg 1$ when the flow reaches a statistically stationary state to avoid initial transients. Then we start to analyse the flow for $T_{s}\leq t\leq T_{s}+\tilde{t}$. Their values are $T_{s}=5000=2500(L/U)$ and $\tilde{t}$ varying from $200=100(L/U)$ to $10000=5000(L/U)$. In Table 1, we list numerical parameters for three representative Reynolds numbers. The convergence study is carried out for $Re=15000$ and $Re=16000$ by increasing the collocation points to $129^{2}$ or $257^{2}$ and halving the time step $\Delta t$ to verify that the results do not change.
As a global fluid mechanical quantity, we consider the kinetic energy,
@@
and use it as an indicator of the flow state as in Shen (1991). We show the variation of the kinetic energy in Figure 2 and its power spectral density (PSD) in Figure 3. The PSD for $Re=14000$ indicates that the flow is periodic. For $Re=15500$, Shen (1991) concluded that the flow is quasi-periodic. The PSD shown in Figure 3(b) is consistent with the conclusion. For $Re=16000$, the PSD becomes broad and the fine structures seen at $Re=15500$ are swallowed by the broad parts. This indicates that the flow is chaotic at $Re=16000$.
ReRe collocation points $\Delta t$ State Shen (1991) State (present)